i need to plot the graphs of equations and use the X and Y intercept method, and a check point. Using the X value when solving for the chechpoint.
4x-3y=12
checkpoint is: 1
i got for the x intercept (3,0)
i got for the y intercept (0, -4)
but i am having troubles with the checkpoint (1,?)

Answer by stanbon(75887)

i need to plot the graphs of equations and use the X and Y intercept method, and a check point. Using the X value when solving for the chechpoint.
4x-3y=12
checkpoint is: 1
i got for the x intercept (3,0)
i got for the y intercept (0, -4)
but i am having troubles with the checkpoint (1,?)
----------------
Let x = 1 and solve for "y":
4*1-3y = 12
-3y = 8
y = -8/3
====
Checkpoint(1, -8/3)
======================
Cheers,
Stan H.